# For and some appropriate functions in [2] the authors considered the system of equations

### v1

@@ 3. Preparatory results
Let , and be as in Proposition 3.4. Let be the solution of the backward problem (1.2). Then from (3.5) and Proposition 3.4 (i), we observe that the inverse problem of recovering the initial value from the knowledge of and is same as solving the operator equation
@@
From Theorem 2.1 it follows that and thus it follows that is a compact operator of infinite rank. Therefore, solving (3.9) is an ill-posed problem. Moreover, if
@@ 4. Quasi-boundary value method and its modification
We have seen in the preceding section that the considered inverse problem of retrieving the initial value from the knowledge of final value and the source term, is an ill-posed problem in the sense that a small perturbations in the data and may lead to a large deviation in the sought solution. But, in practice, we shall have knowledge of the data that are corrupted by some noise. Thus, we have to employ some regularization method to obtain stable approximations. In this section we shall employ the quasi-boundary value method, its modifications and obtain error estimates for apriori as well as aposteriori parameter choice strategies.
@@ 4.1. Error estimates with aposteriori parameter choice
From Proposition 3.4 (ii), we first note that

### v2

@@ 3. Preparatory results
Let , and be as in Proposition 3.4. Let be the solution of the backward problem (1.2). Then from (3.5) and Proposition 3.4 (i), we observe that the inverse problem of recovering the initial value from the knowledge of and is same as solving the linear operator equation
@@
From Theorem 2.1 it follows that and thus it follows that is a compact linear operator of infinite rank. Therefore, solving (3.9) is an ill-posed problem. Moreover, if
@@ 4. Quasi-boundary value method and its modification
We have seen in the preceding section that the considered inverse problem of retrieving the initial value from the knowledge of final value and the source term, is an ill-posed problem in the sense that a small perturbations in the data and may lead to a large deviation in the sought solution. But, in practice, we shall have knowledge of the data that are corrupted by some noise. Thus, we have to employ some regularization method to obtain stable approximations. In this section we shall employ the quasi-boundary value method and its modifications, and obtain error estimates for apriori as well as aposteriori parameter choice strategies.
@@ 4.1. Error estimates with aposteriori parameter choice
From Proposition 3.4 (ii), we first note that for